# MapGPT: Map-Guided Prompting for Unified Vision-and-Language Navigation

### 14 Jan 2024 v1

@@ Abstract
Embodied agents equipped with GPT as their brain have exhibited extraordinary thinking and decision-making abilities across various tasks. However, existing zero-shot agents for vision-and-language navigation (VLN) only prompt the GPT to handle excessive environmental information and select potential locations within localized environments, without constructing an effective “global-view” (e.g., a commonly-used map) for the agent to understand the overall environment. In this work, we present a novel map-guided GPT-based path-planning agent, dubbed MapGPT, for the zero-shot VLN task. Specifically, we convert a topological map constructed online into prompts to encourage map-guided global exploration, and require the agent to explicitly output and update multi-step path planning to avoid getting stuck in local exploration. Extensive experiments demonstrate that our MapGPT is effective, achieving impressive performance on both the R2R and REVERIE datasets (38.8% and 28.4% success rate, respectively) and showcasing the newly emerged global thinking and path planning capabilities of the GPT model. Unlike previous VLN agents, which require separate parameters fine-tuning or specific prompt design to accommodate various instruction styles across different datasets, our MapGPT is more unified as it can adapt to different instruction styles seamlessly, which is the first of its kind in this field.
@@ 1 Introduction
Large language models (LLMs) [27, 8, 41, 42, 14, 6] have demonstrated strong performance in various domains. As the most powerful LLMs, the GPT series models [8, 27, 28, 29] can even serve as the brain of embodied agents [21, 43, 26, 1], enabling them to engage in explicit thinking and decision-making process. Moreover, these GPT-based agents are typically zero-shot or few-shot, eliminating the burdensome tasks of data annotation and model training, and they also demonstrate remarkable generalization abilities.
@@
Recently, LLMs have also been adopted in vision-and-language navigation (VLN), where agents are given human instructions that require them to visually perceive and navigate an indoor environment. Previous learning-based VLN methods [5, 16, 31, 12, 33, 45] relied on the training on large-scale domain-specific data with expert instruction annotations to execute navigation tasks. To address their reliance on training data and poor generalization, some GPT-based zero-shot agents [48, 24] with explainable decision-making abilities have been proposed to tackle the R2R dataset [5]. These methods focused on translating visual observations into textual prompts and required GPT-4 to act as an agent to select the correct position or direction.
@@
However, such zero-shot VLN agents face several challenges. Firstly, transforming all visual observations into text may incorporate excessive useless and distractive information, thereby affecting the performance of the agent. Secondly, to enhance performance, these methods employ multiple additional experts to handle various subtasks, such as summarizing the history and fusing repeated expert predictions, which significantly increases the consumption of time and computational resources. Most importantly, these agents make decisions based solely on the observations of the local environment. This often makes them stuck in local exploration. As shown in Figure 1 (left), given only the local action space, when an agent realizes that it has engaged in an erroneous exploration, it can only continue to explore surrounding environment aimlessly.
We believe that the fundamental reason behind the struggling performance of these zero-shot agents lies in the absence of “global-view” to assist the agent in understanding the spatial structure of the real world, thereby hindering an efficient global exploration. We therefore introduce a simple yet efficient prompt system with only one navigation expert that prioritizes the processing of key node information (i.e., navigable viewpoints) in the environment. We further propose a map-guided prompting method for the GPT model to construct a zero-shot agent, named MapGPT, which aims to provide a unified solution for the VLN task. Specifically, we convert an online constructed topological map into textual prompts and feed them to the GPT model to assist in building and understanding the real-world navigation environment. Leveraging this map-guided prompt, our agent can engage in a more rational global exploration, as shown in Figure 1 (right). Interestingly, by applying the map-guided prompt, we have also uncovered a previously unknown capability of the GPT model, namely multi-step path planning. We require the agent to explicitly perform multi-step path planning and feed the planning as a reference input to the subsequent step, allowing iterative updates to the planning. As a result, the agent can perform explainable global planning to determine the exploration goals for several upcoming steps, or discover an appropriate backtracking path to a prior node for re-exploration if the ongoing exploration encounters obstacles.
We demonstrate impressive performance of MapGPT on the R2R [5] and REVERIE [32] datasets with different styles of instructions. Benefiting from the map-guided prompting and planning, our method outperforms all the existing zero-shot agents on the R2R significantly (around 4% improvement in success rate). In the unexplored zero-shot setting of the REVERIE, our agent exhibits enhanced competitiveness (28.4% success rate), surpassing even some learning-based methods trained on the REVERIE and only being inferior to certain approaches that incorporate additional information, such as precise GPS coordinates.
Benefiting from the flexible single-expert system, as well as the general map-guided prompts for navigation tasks, our MapGPT is more unified. With only some necessary prompt modifications, it can adapt to varying styles of instructions, without the need for extra decomposition of step-by-step instructions (not applicable to high-level instructions) as required by zero-shot DiscussNav [24], or separate parameter fine-tuning on individual datasets as in previous learning-based methods. Therefore, our MapGPT is the first unified approach in the VLN field, allowing effortless adaptation to different downstream tasks in this domain. Besides, MapGPT is applicable to multiple LLMs (e.g., GPT-4 and GPT-3.5) and we also develop the first VLN agent based on the recently released GPT-4V.
@@
  • We propose a novel map-guided path-planning agent, MapGPT, which converts a map constructed online into prompts for GPT, encouraging global exploration.
  • MapGPT achieves impressive performance on the R2R and REVERIE datasets, surpassing existing zero-shot agents. Besides, it is the first unified agent capable of adapting to varying instruction styles effortlessly.
  • By utilizing our novel map-guided prompting method, we uncover a previously unknown ability of the GPT model, namely multi-step path-planning in the real world.
@@ 2 Related Work
Vision-and-Language Navigation (VLN). As a representative multi-modal embodied AI task, VLN requires an agent to combine human instructions and visual observations to navigate and locate targets in real-world scenes. Previous learning-based approaches [44, 25, 15, 31] proposed various model architectures and trained their models on domain-specific datasets. Besides, pretrained models [20, 11, 12, 33, 18, 2, 34, 45, 30] have also been widely applied to this task to produce better multi-modal representations. Recently, to address the reliance on domain-specific data and explore a general agent in diverse scenarios, some zero-shot agents based on GPT [48, 24] have recently been proposed. However, they suffer from several limitations. For example, NavGPT [48] has limited performance and relies on a complex prompt system to process various environmental information. DiscussNav [24] introduces a sequential multi-experts system to discuss environmental information and fuses five repeated predictions to improve performance. Both of these two methods employ multiple experts with different roles to handle sub-tasks such as summarizing observations and decomposing instructions. These designs limit the agent’s capability to only address step-by-step instructions in the R2R dataset [5], and have not been validated on other styles of instructions (e.g., REVERIE [32]). Besides, these agents are limited to local exploration as they can only reason and make decisions within adjacent navigable points. In this paper, we propose a unified and efficient map-guided prompting for global exploaration and multi-step path planning, achieving impressive performance on both R2R and REVERIE.
@@
Large Language Models (LLMs). LLMs [27, 28, 29, 46, 8, 13, 41, 42, 6] have demonstrated remarkable capabilities in multiple domains. Recently, LLM-based agents [21, 43, 26, 1, 47, 37, 30, 7, 36] have also attracted significant interest of the AI community. For example, VoxPoser [21] utilizes LLM and vision-language models to extract affordances and constraints, which enables motion planners to generate trajectories for manipulation tasks. SayCan [1] integrates language models and affordance functions to generate practical strategies for accomplishing the household task. LangNav [30] employs LLMs for navigation, but it merely utilizes GPT-4 [27] to synthesize some data and performs fine-tuning using Llama2 [42] as the backbone, rather than directly employing LLM as a zero-shot agent. In fact, the application of LLM-based zero-shot agents in navigation tasks is still limited, and how to prompt LLM to activate some crucial abilities required by navigation task have yet to be explored. In this paper, we propose a map-guided prompting method and uncover a previously unknown multi-step path-planning ability of the GPT models.
@@
Maps for Navigation. Maps used for navigation tasks can be primarily categorized into two types, i.e., metric maps and topological maps. Employing SLAM [17] for constructing metric maps [9, 40] is widely used in navigation. However, this type of approach requires a trade-off between map size and computational efficiency, which affects navigation performance. To address this limitation, graph-based topological maps [10, 12, 2] have been proposed for pre-exploring environment or enabling global exploration, such as backtracking to previously visited nodes. However, these methods are all designed for model learning. It remains unexplored how to construct a map prompt and leverage the powerful capabilities of LLMs for zero-shot reasoning and planning based on the map.
@@
Figure 2: The framework of our proposed MapGPT. Our basic system consists of two types of prompts, namely task description and fundamental inputs for VLN. We further introduce a map-guided prompting method that converts an online-constructed topological map into prompts, activating the agent’s global exploration. Finally, we require the agent to explicitly update multi-step navigation planning for each step based on the previous one, to enhance navigation performance.
@@ 3 Method
In this section, we first introduce our newly designed prompt system for the VLN task (Section 3.1). We then describe the details of our novel map-guided prompting method to establish the agent’s understanding of the global spatial structure of the environment (Section 3.2), which encourages the agent to explicitly generate and update its multi-step planning (Section 3.3).
3.1 Unified Single Expert Prompt System
Designing appropriate prompts for LLMs to activate their various abilities is of critical importance. However, prior NavGPT [48] and DiscussNav [24] gathered excessive noisy environmental information from all the views, which may interfere with the GPT’s decision-making. Besides, they relied on complex multi-expert systems, where GPT [27] played different roles to achieve various functions, such as instruction parsing, summarizing observations and history, etc. This undoubtedly increases resource consumption and makes it challenging to achieve seamless collaboration among multiple experts with different backgrounds. While a single expert may have a comprehensive understanding of complex task background and is capable of handling various sub-tasks uniformly.
As a comparison, our designed single expert prompt system has several features. (1) Utilizing one single navigation expert, we eliminate the need for a separate design of additional historical summary expert and progress estimation expert based on GPT models. (2) It is unified and flexible. We do not contain any customized design specific to instruction style, and it is convenient to incorporate additional information, such as maps. (3) We prioritize the observations of navigable points, which are combined with the corresponding directions to form a continuous semantics in action space, facilitating comprehension and decision-making.
As shown in Figure 2, we adopt a modular design for our prompt system. At each step, we collect various fundamental inputs for the agent, including instruction $I$, history $H_{t}$, observation $O_{t}$, and action space $A_{t}$. The meaning of these inputs, as well as the requirements for output, are clearly pre-defined in the task description $D$. We utilize a prompt manager $PM$ to organize these prompt inputs which are then fed into the large language model $LLM$ to generate the current thought $T_{t}$ and select a specific action $a_{t,i}\in A_{t}$. The pipeline can be formulated as
@@ 3.1.1 Task Description
As shown in Figure 2, our prompts for task description $D$ consist of three main parts. Firstly, we set the task background by telling GPT that “You are an embodied robot that navigates in the real world …”. Secondly, we provide some definitions regarding the inputs at each step. For example, “‘Action options’ are some actions that you can take at this step”. Thirdly, we propose some basic requirements regarding how the agent should accomplish this task, such as some format requirements and an emphasis of thoughtful consideration like “think about what to do next and why …”.
@@ 3.1.2 Fundamental Inputs
Our agent is equipped with an RGB camera to capture $M$ images of the environment. Unlike NavGPT [48] that gathers excessive noisy environmental information, we utilize only the views that contain navigable points, and translate these observations into scene and object descriptions to prompt GPT to perceive the real world. The only exception is the stopping condition in the REVERIE dataset, which requires finding certain objects that may not be located in navigable directions, such as a painting hanging on a wall. Therefore, we utilize the bounding boxes provided by the REVERIE to extract an additional object list of all the views, named “Surroundings”, to enable the agent to determine whether to stop. Since we do not specify the directions of these surrounding objects, they cannot be used for navigation and are only utilized for stopping the agent in the REVERIE dataset.
To translate the visual observations into text descriptions, we follow NavGPT [48] and utilize off-the-shelf vision models BLIP-2 [23] and Faster R-CNN [35, 4] to acquire scene description and object detection respectively. Finally, we formulate these observations that have been converted into text as $O_{t}=\{o_{t,i}\}_{i=1}^{N}$, where $N$ is the number of navigable viewpoints at step $t$. Each $o_{t,i}$ represents the observation towards a specific navigable point and is formulated using the prompt template:
“<{scene}>, which also includes <{objects}>”.
@@
NavGPT allows the agent to directly select a viewpoint and DiscussNav enables the model to predict a direction. However, we incorporate both directional phrases (e.g., turn left to) and corresponding observation of this direction into the action space for GPT to select from, making the decision-making process more intuitive.
@@
{label} {directional_phrase} {$o_{t,i}$.
We have defined six directional concepts, namely “go forward to”, “turn left to”, “turn right to”, “turn around to”, “go up to”, and “go down to”, according to the directions of navigable viewpoints. For instance, $a_{t,1}$ could be “B. turn around to <a room with blue walls>, which also includes <bed, curtain, picture>”. At each step $t$, the agent only needs to choose one option $a_{t,i}\in A_{t}$. For the output format, we require the agent to simply provide a single option label, such as ”Action: B.”, which is easy to parse.
@@
We directly record all previous actions $a_{0}\sim a_{t-1}$ for history. The following prompt template is utilized for appending the actions into $H_{t}$:
“step 0: {$a_{0}^{\ast}$}, step 1: {$a_{1}^{\ast}$}, …, step t-1: {$a_{t-1}^{\ast}$},

### 25 Feb 2024 v2

# MapGPT: Map-Guided Prompting with Adaptive Path Planning for Vision-and-Language Navigation

@@ Abstract
Embodied agents equipped with GPT as their brain have exhibited extraordinary decision-making and generalization abilities across various tasks. However, existing zero-shot agents for vision-and-language navigation (VLN) only prompt the GPT-4 to select potential locations within localized environments, without constructing an effective “global-view” for the agent to understand the overall environment. In this work, we present a novel map-guided GPT-based agent, dubbed MapGPT, which introduces an online linguistic-formed map to encourage the global exploration. Specifically, we build an online map and incorporate it into the prompts that include node information and topological relationships, to help GPT understand the spatial environment. Benefiting from this design, we further propose an adaptive planning mechanism to assist the agent in performing multi-step path planning based on a map, systematically exploring multiple candidate nodes or sub-goals step by step. Extensive experiments demonstrate that our MapGPT is applicable to both GPT-4 and GPT-4V, achieving state-of-the-art zero-shot performance on the R2R and REVERIE simultaneously ($\sim$10and $\sim$12improvements in SR), and showcasing the newly emerged global thinking and path planning abilities of the GPT.
MapGPT: Map-Guided Prompting with Adaptive Path Planning for Vision-and-Language Navigation
[…]
@@ 1 Introduction
Large language models (LLMs) Touvron et al. (2023a, b); Chowdhery et al. (2022); Anil et al. (2023) have demonstrated strong performance in various domains. As the most powerful LLMs, the GPT series models Brown et al. (2020); OpenAI (2023a, b, c) can even serve as the brain of embodied agents Huang et al. (2023); Mu et al. (2023); Ahn et al. (2022), enabling them to engage in explicit thinking and decision-making process. Moreover, these GPT-based agents are typically zero-shot or few-shot, eliminating the burdensome tasks of data annotation and model training, and they also demonstrate remarkable sim-to-real abilities.
@@
Recently, LLMs have also been adopted in vision-and-language navigation (VLN), where agents are given human instructions that require them to visually perceive and navigate an indoor environment. Previous learning-based VLN methods Anderson et al. (2018b); Fried et al. (2018); Qi et al. (2020a); Chen et al. (2022); Qiao et al. (2022); Wang et al. (2023) relied on the training on large-scale domain-specific data with expert instruction annotations to execute navigation tasks. To address their reliance on training data and potential sim-to-real gap, some GPT-based zero-shot agents Zhou et al. (2023); Long et al. (2023) with explainable decision-making abilities have been proposed. These methods translated visual observations into textual prompts and required GPT-4 to act as an agent to select the correct position or direction. Besides, they also employed multiple additional experts to handle various subtasks in text form, such as summarizing the history and fusing repeated expert predictions.
@@
However, such zero-shot VLN agents face some challenges. Firstly, these methods are developed for a multi-expert pure language system based on GPT-4, necessitating the conversion of visual observations into text and multiple rounds of text summarizations, which inevitably leads to information loss and hinders the application to multimodal LLMs (e.g., GPT-4V). More importantly, these agents make decisions based solely on the observations of the local environment. As shown in Figure 1 (left), given only the local action space, when an agent realizes that it has engaged in an erroneous exploration, it can only continue to explore surrounding environment aimlessly.
In this paper, we propose MapGPT which contains a topological map in the linguistic form to assist in global exploration and adaptive path planning. We first develop a simple yet efficient prompt system with only one navigation expert that can be applied to both GPT-4 and multimodal GPT-4V flexibly. To encourage global exploration, we propose a map-guided prompting method for the GPT model, to build a “global-view” for the agent. Specifically, for an online constructed map, we have discarded the precise GPS coordinates that are difficult for GPT to understand, while preserving the topological relationships between nodes and incorporating them into prompts to assist in understanding the navigation environment, as shown in Figure 1 (right). Given this tailored map, we further propose an adaptive planning mechanism to activate GPT’s multi-step path planning ability. Instead of documenting the thinking process of each step as in previous works, MapGPT generates a multi-step planning similar to a human work plan, and updates it iteratively to achieve strategic exploration of potential objectives. Benefiting from this, the agent can adaptively perform path planning based on the map, systematically explore multiple candidate nodes or sub-goals step by step, and backtrack to a specific node for re-exploration when necessary.
We conduct experiments on two popular VLN benchmarks, namely R2R Anderson et al. (2018b) and REVERIE Qi et al. (2020b), which contain step-by-step and high-level instructions respectively. Experimental results show the superiority of MapGPT over existing zero-shot VLN agents. Especially in REVERIE, MapGPT exhibits enhanced competitiveness (31.6% SR), surpassing even some learning-based methods trained on REVERIE. Extensive ablation studies reveal the advantage of our introduced map and adaptive path planning mechanism in encouraging systematic exploration to improve navigation.
@@
  • We propose a novel map-guided prompting method, which introduces an online linguistic-formed map including node information and topological relationships to encourage GPT’s global exploration.
  • An adaptive planning mechanism is utilized to activate GPT’s multi-step path-planning ability, enabling systematic exploration of multiple potential objectives.
  • MapGPT can be applied to both GPT-4 and GPT-4V and is more unified as it can adapt to varying instruction styles effortlessly, achieving state-of-the-art zero-shot performance on both the R2R and REVERIE datasets.
@@ 2 Related Work
Vision-and-Language Navigation (VLN) As a representative multi-modal embodied AI task, VLN requires an agent to combine human instructions and visual observations to navigate and locate targets in real-world scenes. Previous learning-based approaches Wang et al. (2019); Ma et al. (2019); Deng et al. (2020); Qi et al. (2020a) proposed various model architectures and trained their models on domain-specific datasets. Besides, pretrained models Hong et al. (2021); Chen et al. (2021b, 2022); Qiao et al. (2022); Guhur et al. (2021); An et al. (2022); Qiao et al. (2023); Wang et al. (2023); Pan et al. (2023) have been widely applied to produce better multi-modal representations. Recently, to address the reliance on domain-specific data and the possible sim-to-real gap, some zero-shot agents based on GPT Zhou et al. (2023); Long et al. (2023) have been proposed. However, they suffer from several limitations. For example, NavGPT Zhou et al. (2023) has limited performance and relies on a two-stage language-only system. DiscussNav Long et al. (2023) introduces a sequential multi-experts system to discuss and summarize various information and fuses five repeated predictions to improve performance. Some of their designs limit the agent’s capability to only address step-by-step instructions in the R2R dataset Anderson et al. (2018b), and have not been validated on other styles of instructions (e.g., REVERIE Qi et al. (2020b)). Besides, these agents are limited to local exploration as they can only reason and make decisions within adjacent navigable points. In this paper, we propose a map-guided prompting method with adaptive path planning for global exploration, achieving impressive performance on both R2R and REVERIE.
@@
Large Language Models (LLMs) LLMs OpenAI (2023a, b, c); Chiang et al. (2023); Touvron et al. (2023a, b); Anil et al. (2023) have demonstrated remarkable capabilities in multiple domains. Recently, LLM-based agents Huang et al. (2023); Mu et al. (2023); Ahn et al. (2022); Pan et al. (2023); Brohan et al. (2023); Schumann et al. (2023); Hu et al. (2023) have also attracted significant interest of the community. For example, VoxPoser Huang et al. (2023) utilizes LLM and vision-language models to extract affordances and constraints, which enables motion planners to generate trajectories for manipulation. LangNav Pan et al. (2023) employs LLMs for navigation, but it merely utilizes GPT-4 OpenAI (2023a) to synthesize some data and performs fine-tuning using Llama2 Touvron et al. (2023b) as the backbone, rather than directly employing LLM as a zero-shot agent. In fact, the application of LLM-based zero-shot agents in navigation tasks is still limited, and how to prompt LLMs (including multimodal GPT-4V) to activate global thinking and planning abilities required by navigation task have yet to be explored Yang et al. (2023).
@@
Maps for Navigation Maps used for navigation tasks can be primarily categorized into two types, i.e., metric maps and topological maps. Employing SLAM Fuentes-Pacheco et al. (2015) for constructing metric maps Chaplot et al. (2020); Thrun (1998) is widely used in navigation. However, this type of approach requires a trade-off between map size and computational efficiency, which affects navigation performance. To address this limitation, graph-based topological maps Chen et al. (2021a, 2022); An et al. (2022) have been proposed for pre-exploring environment or enabling global exploration, such as backtracking to previously visited nodes. However, these methods are all designed for model learning. It remains unexplored how to build a map with prompts and leverage the powerful capabilities of LLMs for zero-shot reasoning and planning based on the map.
@@
Figure 2: Our basic system consists of two types of prompts, namely task description and fundamental inputs. We introduce a map-guided prompting method that builds an online-constructed topological map into prompts, activating the agent’s global exploration. We further propose an adaptive mechanism to perform multi-step path planning on this map, systematically exploring candidate nodes or sub-goals. Note that vision models are optional, and viewpoint information can be represented using either the image or textual description of the observations.
@@ 3 Method
In this section, we introduce our newly designed prompt system (Sec. 3.1), the details of our map-guided prompting method to help the agent understand global environment (Sec. 3.2), and the novel adaptive mechanism that encourages the agent to make multi-step path planning (Sec. 3.3).
3.1 Single Expert Prompt System
Previous works such as NavGPT Zhou et al. (2023) and DiscussNav Long et al. (2023) are two-stage systems. They first gather visual observations from all the views and translate them into textual descriptions which are then fed into language-only GPT-4 OpenAI (2023a) for decision-making. Besides, they rely on complex multi-expert designs, where GPT plays different roles to achieve various functions, such as instruction parsing, summarizing textual descriptions and history, etc. However, these intricate designs are only geared towards text processing, which limits their research value. Powerful multimodal large models, such as GPT-4V OpenAI (2023b, c), can directly serve as the agent’s brain to process multimodal information and make decisions Yang et al. (2023).
As a comparison, our designed single expert prompt system has several features. (1) We eliminate the need for a separate design of an additional historical summary expert or instruction decomposing expert based on GPT models, which makes it convenient to incorporate both visual/textual inputs and additional information, such as maps. (2) Our navigation expert can utilize GPT-4V to make decisions directly based on visual observations in one stage. It can also take text descriptions as inputs and flexibly apply them in the two-stage GPT-4 system. (3) Simple yet efficient. In the R2R dataset, our two-stage system requires an average of 672 input tokens and 115 output tokens per step. In comparison, NavGPT utilizes three GPT experts, resulting in an average cost of 2,465 input tokens and 317 output tokens per step.
As shown in Figure 2, we collect various fundamental inputs for the agent, including instruction $I$, history $H_{t}$, observation $O_{t}$, and action space $A_{t}$. The meaning of these inputs, as well as the requirements for output, are clearly pre-defined in the task description $D$. We utilize a prompt manager $PM$ to organize these prompt inputs which are then fed into the large language model $LLM$ to generate the current thought $T_{t}$ and select a specific action $a_{t,i}\in A_{t}$. The pipeline can be formulated as
@@ 3.1.1 Task Description
As shown in Figure 2, our prompts for task description $D$ consist of three main parts. We set the task background for GPT, provide some definitions regarding the inputs at each step, and propose some basic requirements regarding how the agent should accomplish this task.
@@ 3.1.2 Fundamental Inputs
Our agent is equipped with an RGB camera to capture $M$ images of the environment. Unlike NavGPT Zhou et al. (2023) that gathers excessive environmental information, we prioritize the observations of navigable points, which can be formulated as $O_{t}=\{o_{t,i}\}_{i=1}^{N}$, where $N$ is the number of navigable viewpoints at step $t$. Each $o_{t,i}$ represents the observation towards a specific navigable point. These observations are original images for a one-stage system based on GPT-4V and can also be replaced with caption and detection results in a two-stage system, where we follow NavGPT Zhou et al. (2023) and utilize off-the-shelf vision models BLIP-2 Li et al. (2023) and Faster R-CNN Ren et al. (2015); Anderson et al. (2018a) to acquire scene description and object detection respectively. NavGPT also utilizes the bounding boxes provided by the REVERIE dataset to extract an additional object list of all the views, which is then used for the R2R experiment. We also utilize this object list and name it the “Surroundings”. However, we do not utilize this additional information in R2R. We only utilize it to enable the agent to determine whether to stop in the REVERIE dataset since we do not specify the directions of these surrounding objects.
@@
NavGPT allows the agent to directly select a viewpoint and DiscussNav enables the model to predict a direction. In this paper, we incorporate both directional phrases (e.g., turn left to) and the corresponding observations of each direction into the action space for GPT to select from, making the decision-making process more intuitive.
@@
{label} {direction} {$o_{t,i}$.
At each step $t$, the agent only needs to choose one option $a_{t,i}\in A_{t}$. For the output format, we require the agent to simply provide a single option label, such as "Action: B".
@@
We record all previous actions $a_{0}\sim a_{t-1}$ for history. The following prompt template is utilized for appending the actions into $H_{t}$:
“step 0: {$a_{0}^{\ast}$}, …, step t-1: {$a_{t-1}^{\ast}$},